calculate the capacitance of a capacitor if it stores a charge of 2.3x10-4 when delivered with a voltage of 4x103V



1
Expert's answer
2022-04-21T13:01:13-0400

Given:

"q=2.3*10^{-4}\:\rm C"

"V=4*10^{3}\:\rm V"


The capacitance

"C=\frac{q}{V}=\frac{2.3*10^{-4}}{4*10^3}=5.75*10^{-8}\:\rm F"